OpenStudy (anonymous):

Hi! Please HELP! I am not looking for answers! I need to solve for x: 62.10=(20+0.5x)+0.15(20+0.5x)

OpenStudy (anonymous):

This what I know so far. I need to times 0.15 by 20 and by 0.5. but I'm confused on where I go from there. Please help me solve this!

OpenStudy (anonymous):

\[62.10=(20+0.5x)+0.15(20+0.5x)\]

OpenStudy (kainui):

So after you've done that notice that you can remove all the parenthesis. Does that make it simpler to look at? From there you can combine things. For instance if you have 2 apples + 3 apples +7 oranges you can combine the apple terms together to get 5 apples + 7 oranges. Similarly you'll have something like 100x+34+20x+7 you can combine terms in this little example to get 120x+41.

OpenStudy (anonymous):

Oops, typed that out wrong: 62.10=20+0.5+3+0.075?

OpenStudy (texaschic101):

62.10 = (20 + 0.5x) + 0.15(20 + 0.5x) 62.10 = 20 + 0.5x + 3 + 0.075x 62.10 = 23 + .575x 62.10 - 23 = .575x 39.10 = .575x 39.10/.575 = x 68 = x check.. 62.10 = 20 + .5(68) + .15(20 + .5(68) 62.10 = 20 + 34 + .15(20 + 34) 62.10 = 54 + .15(54) 62.10 = 54 + 8.10 62.10 = 62.10 (correct) x = 68

OpenStudy (anonymous):

Oh, okay, I get it! You combine the terms!

OpenStudy (anonymous):

Thank you to both @texaschic101 & @Kainui!! :)